# Efficient first-principles inverse design of nanolasers

### v2

@@ Appendix B Lasing mode amplitude in the weak pumping regime
In this appendix, we detail the calculation of the lasing amplitude in the high-$Q$, small amplitude, and small pumping strength limit. It is possible to expand Eq. (7) for small amplitudes ($|a_{0}|^{2}\ll 1$), including the zeroth and the first-order terms. This gives
$\Delta\varepsilon_{\Im}(\mathbf{r})\approx D_{0}(\mathbf{r})d\left(1-|a_{0}|^{2}\,\frac{e_{c}^{-2}|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}}{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\right)\,.$ (B1)
@@
$|a_{0}|^{2}\approx\frac{\Delta d}{d_{\text{thresh}}}\frac{{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{e_{c}^{-2}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{4}\,\mathop{}\!\mathrm{d}\Omega}\,,$ (B2)
where $\Delta d$ denotes difference of the pumping strength $d$ with the threshold. (At large amplitudes, many additional nonlinear effects arise, including competition from additional lasing modes, eventually requiring a full nonlinear solver [39].)
@@ Appendix C Lasing threshold and FOM with steady-state diffusion
In this appendix, we revisit the derivations of the laser properties (e.g., threshold, lasing mode intensity) and FOM, including steady-state diffusion effects through the C-SALT formalism [4]. We introduce diffusion in the gain model in Eq. (16). By rewriting the lasing mode as a function of the passive cavity mode, we find
$\Delta\varepsilon_{\Im}(\mathbf{r})=\mathbb{S}\left(\mathbb{I}+|a_{0}|^{2}\,\frac{e_{c}^{-2}|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}}{\int_{\Omega}\varepsilon(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\mathbb{S}\right)^{-1}[D_{0}](\mathbf{r})d\,.$ (C1)
Similar to the case without diffusion, close to the lasing threshold, we assume a small amplitude $(|a_{0}|^{2}\ll 1)$, so we can expand the expression in a Taylor series around $|a_{0}|^{2}=0$. For zero amplitude the change in permittivity is $\Delta\varepsilon_{\Im}(\mathbf{r})=\mathbb{S}[D_{0}](\mathbf{r})d\,.$ Using this expression together with the lasing condition in Eq. (8), gives an updated version of the threshold pumping strength that accounts for diffusion effects, resulting in Eq. (17).
@@
$\Delta\varepsilon_{\Im}(\mathbf{r})\approx\mathbb{S}\left(\mathbb{I}-|a_{0}|^{2}\,\frac{e_{c}^{-2}|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}}{\int_{\Omega}\varepsilon(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\mathbb{S}\right)[D_{0}](\mathbf{r})d\,.$ (C2)
@@
$|a_{0}|^{2}\approx\frac{\Delta d}{d_{\text{thresh}}}\frac{{\int_{\Omega}\varepsilon(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\int_{\Omega}\mathbb{S}[D_{0}](\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{e_{c}^{-2}\int_{\Omega}\mathbb{S}\left[|\mathbf{E}_{\mathrm{m}}|^{2}\,\mathbb{S}[D_{0}]\right](\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,.$ (C3)
@@
Appendix D Simulation and optimization parameters
In this section we detail the parameters used in the simulation and optimization of the nanolaser devices. The numerical model is discretized using the finite-element method [55], with first-order Nedelec elements [55] for the frequency-domain electromagnetic problem and linear Lagrange elements for the steady-state exponentially damped diffusion problems.
The simulations are performed in COMSOL Multiphysics (version 6.2) [73] and the optimization is solved using the Globally Convergent Method of Moving Asymptotes (GCMMA) [74]. All computations are executed on the DTU Computing Center HPC cluster [75].
@@ Appendix D Simulation and optimization parameters
Two dimensional systems: In the 2d computational results (Sec. 3.1, 3.2) we consider a geometry based on a rectangular simulation domain with length $l=13.95\,\upmu\text{m}$ (along the $x$ direction) and width $w=4.55\,\upmu\text{m}$ (along the $y$ direction), where the rectangular design domain ($\Omega$) is centered at position $\mathbf{r}_{\text{0}}=(1.1625\,\upmu\text{m}$, $0\,\upmu\text{m}$), and has side of length of $l_{\Omega}=2\lambda=3.1\,\upmu\text{m}$. The design domain is connected to a waveguide with width $w_{\text{wg}}=500$ nm, which is excited at position $\mathbf{r}_{\text{src}}=(-1\,\upmu\text{m}$, $0\,\upmu\text{m}$).
Three dimensional systems: In the 3d computational results (Sec. 3.3), we extend the previous geometry by extruding the two-dimensional design domain ($\Omega$) and waveguide out-of-plane (along the $z$-direction) to a height of $h_{\Omega}=220$ nm, and placing the resulting structure on a silica substrate. The design domain is at the center of a simulation domain with total height of $h=2\lambda=3.1\,\upmu\text{m}$.
@@
The main optimization parameter choices are the threshold value $\eta=0.5$ and filter radius $r_{\text{f}}=100$ nm. The geometric length scale constraints [51] are enforced using the parameters $c=39$, and $\eta_{e}=0.75$ and $\eta_{d}=0.25$, as to ensure a lengthscale $\sim r_{\mathrm{f}}/2=50$ nm [52]. The rest of the parameters, including the length scale error ($\epsilon$) needed to enforce the length scale constraints  [51], are chosen as part of a continuation scheme, which is essential to enable the use of gradient-based methods and ensures a well-performing final binary design. In the continuation scheme we vary the simulation parameters as a function of the iteration number to find a binary and well performing final design. We summarize those parameter choices in Tab. 1.
@@
Continuation step 00 1 2 3 4 5 6 7 8
Iteration 00 100 200 300 400 500 600 700 800

### v3

@@ Appendix B Lasing mode amplitude in the weak pumping regime
Following a similar derivation to the single mode lasing amplitude in App. A, in this appendix we detail the calculation of the lasing amplitude in the high-$Q$, small amplitude, and small pumping strength limit. It is possible to expand Eq. (7) for small amplitudes ($|a_{1}|^{2}\ll 1$), including the zeroth and the first-order terms. This gives
$\Delta\varepsilon_{\Im}(\mathbf{r})\approx D_{0}(\mathbf{r})d\left(1-|a_{1}|^{2}\,\frac{e_{c}^{-2}|\mathbf{e}_{1}(\mathbf{r})|^{2}}{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\right)\,.$ (B1)
@@
$|a_{1}|^{2}\approx\frac{\Delta d}{d_{\text{thresh}}}\frac{{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{e_{c}^{-2}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{4}\,\mathop{}\!\mathrm{d}\Omega}\,$, (B2)
where $\Delta d$ denotes difference of the pumping strength $d$ with the threshold. (At large amplitudes, many additional nonlinear effects arise, including competition from additional lasing modes, eventually requiring a full nonlinear solver [EsterhazyLiu14].)
@@ Appendix C Lasing threshold and FOM with steady-state diffusion
In this appendix, we revisit the derivations of the laser properties (e.g., threshold, lasing mode intensity) and FOM, including steady-state diffusion effects through the C-SALT formalism [csalt]. We introduce diffusion in the gain model in Eq. (16). By rewriting the lasing mode as a function of the passive cavity mode, we find
$\Delta\varepsilon_{\Im}(\mathbf{r})=\mathbb{S}\left(\mathbb{I}+|a_{1}|^{2}\,\frac{e_{c}^{-2}|\mathbf{e}_{1}(\mathbf{r})|^{2}}{\int_{\Omega}\varepsilon(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\mathbb{S}\right)^{-1}[D_{0}](\mathbf{r})d\,.$ (C1)
Similar to the case without diffusion, close to the lasing threshold, we assume a small amplitude $(|a_{1}|^{2}\ll 1)$, so we can expand the expression in a Taylor series around $|a_{1}|^{2}=0$. For zero amplitude the change in permittivity is $\Delta\varepsilon_{\Im}(\mathbf{r})=\mathbb{S}[D_{0}](\mathbf{r})d\,.$ Using this expression together with the lasing condition in Eq. (8), gives an updated version of the threshold pumping strength that accounts for diffusion effects, resulting in Eq. (17).
@@
$\Delta\varepsilon_{\Im}(\mathbf{r})\approx\mathbb{S}\left(\mathbb{I}-|a_{1}|^{2}\,\frac{e_{c}^{-2}|\mathbf{e}_{1}(\mathbf{r})|^{2}}{\int_{\Omega}\varepsilon(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\mathbb{S}\right)[D_{0}](\mathbf{r})d\,.$ (C2)
@@
$|a_{1}|^{2}\approx\frac{\Delta d}{d_{\text{thresh}}}\frac{{\int_{\Omega}\varepsilon(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\int_{\Omega}\mathbb{S}[D_{0}](\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{e_{c}^{-2}\int_{\Omega}\mathbb{S}\left[|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathbb{S}[D_{0}]\right](\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,.$ (C3)
@@
Appendix D Validating the SPA-SALT approximations
To validate the SPA-SALT approximation, we will use the derivations in App. A. Primarily, we will evaluate the two key approximations in our SPA-SALT-based optimization framework: that a single cavity mode will lase with a threshold significantly lower than other cavity modes, and that the electric field in the reciprocal problem accurately represents the single lasing mode at threshold and above. As an example, we will evaluate this by using the device optimized in 2d for a gain region with $\sigma=500$ nm (i.e., the cavity design shown in Fig. 3 (d) and Fig. 4).
D.1 Lasing threshold for a second lasing mode
First, we identify the cavity modes closest to the optimized lasing mode. These modes are resonances with a maximum $Q$-factor of $Q\approx 200$ ($\approx$ 1-2 orders of magnitude lower than the lasing mode), which are spectrally separated by just a few nanometers. Given this spectral closeness, we approximate that the closest modes do not experience any line pulling effects ($\gamma_{\perp}\to\infty$, so that $\Gamma_{1}\to 1$ and $\Gamma_{2}\to 1$). By using the expressions for the interacting threshold of a second lasing mode in Eq. (A15) and assuming $\omega_{L_{2}}=\omega_{2}$, we find that the value for the second interacting threshold for the closest modes diverge: $d^{(2)}_{\textrm{int thr}}\to\infty$. This means that within the SPA-SALT approximation a second mode will not reach threshold in the presence of modal interactions. This result thus supports the assumption that our system is dominated by a single lasing mode.
D.2 Linear response validation for high-$Q$ cavities
Next, we validate the error between the linear response of the reciprocal system and the SALT model above the threshold by using the scalar residual measure defined in Eq. (A19). The residual gives a scalar measure of the relative error between the linear solution and the SALT model, which can be used to validate the high-$Q$ approximation (i.e., that the reciprocal field is proportional to the cavity mode).
Figure 7: Validation of the linear response high-$Q$ approximation at and above threshold. The residual as a function of pump strength $d$ for pumping strengths up to $d\approx 5d_{\text{thresh}}$. The lasing threshold is marked in red at $d_{\text{thresh}}=2.01\cdot 10^{-3}$.
To show the validity of this approximation in Fig. 7, we plot the residual (Eq. (A19)) as a function of the scalar pumping strength ($d$). Close to the threshold ($d_{\text{thresh}}$), the residual is very low (a relative error of $\lessapprox 0.4\%$), whereas as the pumping strength increases, the residual and the relative error start to increase. Nevertheless, even for a larger pumping strength of $d\approx 5d_{\text{thresh}}$, the relative error is still smaller than $\lessapprox 1\%$. We note that for even larger pumping strengths, the residual grows as a linear function, and we calculate that for a pumping strength of $d\approx 100d_{\text{thresh}}$ the relative error is still $\approx 5\%$. These results confirm the robustness and validity of the approximation for the inverse-designed high-$Q$ nanolaser cavities.
Appendix E Simulation and optimization parameters
In this section we detail the parameters used in the simulation and optimization of the nanolaser devices. The numerical model is discretized using the finite-element method [jin], with first-order Nedelec elements [jin] for the frequency-domain electromagnetic problem and linear Lagrange elements for the steady-state exponentially damped diffusion problems.
The simulations are performed in COMSOL Multiphysics (version 6.2) [COMSOL] and the optimization is solved using the Globally Convergent Method of Moving Asymptotes (GCMMA) [GCMMA]. All computations are executed on the DTU Computing Center HPC cluster [DTU_DCC_resource].
@@ Appendix E Simulation and optimization parameters
Two dimensional systems: In the 2d computational results (Sec. 3.1, 3.2) we consider a geometry based on a rectangular simulation domain with length $l=13.95\,\upmu\text{m}$ (along the $x$ direction) and width $w=4.55\,\upmu\text{m}$ (along the $y$ direction), where the rectangular design domain ($\Omega$) is centered at position $\mathbf{r}_{\text{0}}=(1.1625\,\upmu\text{m}$, $0\,\upmu\text{m}$), and has side of length of $l_{\Omega}=1.5\lambda=2.325\,\upmu\text{m}$. The design domain is connected to a waveguide with width $w_{\text{wg}}=500$ nm, which is excited at position $\mathbf{r}_{\text{src}}=(-1\,\upmu\text{m}$, $0\,\upmu\text{m}$).
Three dimensional systems: In the 3d computational results (Sec. 3.3), we extend the previous geometry by extruding the two-dimensional design domain ($\Omega$) and waveguide out-of-plane (along the $z$-direction) to a height of $h_{\Omega}=220$ nm, and placing the resulting structure on a silica substrate. The design domain is at the center of a simulation domain with total height of $h=1.5\lambda=2.325\,\upmu\text{m}$.
@@
The main optimization parameter choices are the threshold value $\eta=0.5$ and filter radius $r_{\text{f}}=100$ nm. The geometric length scale constraints [lengthscale] are enforced using the parameters $c=39$, and $\eta_{e}=0.75$ and $\eta_{d}=0.25$, as to ensure a lengthscale $\sim r_{\mathrm{f}}/2=50$ nm [meep]. The rest of the parameters, including the length scale error ($\epsilon$) needed to enforce the length scale constraints  [lengthscale], are chosen as part of a continuation scheme, which is essential to enable the use of gradient-based methods and ensures a well-performing final binary design. In the continuation scheme we vary the simulation parameters as a function of the iteration number to find a binary and well performing final design. We summarize those parameter choices in Tab. 1.
@@
Continuation step 0 1 2 3 4 5 6 7 8
Iteration 0 100 200 300 400 500 600 700 800